# Seyed Ali Farokh

### v1

@@ 4.3 Whisper-PMFA
Building on the successful use of pre-trained ASR models in speaker verification [11, 12], Whisper-PMFA (Partial Multi-Scale Feature Aggregation using Whisper) [30] was recently introduced to leverage the capabilities of Whisper, a large-scale multilingual ASR model based on transformers [34]. Whisper-PMFA adapts Whisper for speaker verification by selectively concatenating frame-level outputs from specific transformer layers, rather than aggregating features from all layers. This approach not only reduces computational overhead but also enhances performance by minimizing the integration of irrelevant information from lower-impact layers.
Inspired by this, we studied the performance of Whisper-PMFA in this challenge. Since Whisper was not trained for the speaker recognition task, we applied both training stages to Whisper-PMFA. Additionally, before the full training stage, we froze the Whisper parameters and fine-tuned the model for five epochs to prevent updating the pre-trained model in the wrong direction due to the random initialization of newly added components.
We submitted two Whisper-PMFA-based systems to this challenge (S5 and S6), with the only difference being the AAM-Softmax margin during the domain adaptation phase, which was set to 0.35 and 0.2 for S5 and S6, respectively.
@@ 4.4 Feature extraction
80-dimensional log Mel filter bank energies with a 25ms window and 10ms frame shift were extracted for our ResNet-based models. Voice activity detection (VAD) was not applied, and all features were mean-normalized. Likewise, 80-dimensional log magnitude Mel spectrograms consistent with the pre-trained Whisper were utilized for training Whisper-PMFA.
@@ 4.5 Backend
Afterward, AS-Norm [35] was used for score normalization, using 1,620 cohorts obtained from speaker-wise averaging of all embeddings in the training subset of the challenge dataset. Top-300 most similar scores were selected to compute the mean and standard deviation for normalization.
@@ 5 Results
Table 3 shows the evaluation results of our single and fusion systems on the development and evaluation subsets of the challenge after applying AS-Norm and rejecting TW trials. The results indicate that the Whisper-PMFA method outperforms the widely used ResNet architecture, conforming to the findings of previous studies on the effectiveness of adapting pre-trained ASR models for speaker verification. However, it can be observed from the results that the ResNets pre-trained on approximately twice the data (VoxCeleb 1&2) can surpass Whisper-PMFA after a simple domain adaptation stage, which highlights the importance of large-scale pre-training in improving the generalization ability of speaker verification models.
@@ 5.1 Ablation study
Table 4: Ablation study
Methods Development
MinDCF0.01 EER(%)
Whisper-PMFA (T1) 0.0234 0.9253
+ Domain adaptation (T2) 0.0177 0.6273
++ AS-Norm 0.0161 0.6126
@@ 6 Conclusion
In this report, we described our system for Task 1 of the Iranian division of the TdSV challenge 2024. Our phrase verification system was a classifier built on top of XLSR, and our speaker verification models were based on Whisper-PMFA and pre-trained ResNets. The experimental results show that, in this task, the performance of Whisper-PMFA is inferior to a pre-trained ResNet293 with domain adaptation, even though it can outperform the ResNet architecture with random initialization and the same training data. The results also demonstrate the possibility of developing competitive text-dependent speaker verification systems using independent pre-trained components. Finally, our best system was a fusion of multiple ResNet- and Whisper-PMFA-based models, which achieved a MinDCF of 0.0358 on the evaluation subset and won the challenge.

### v2

@@ 4.3 Whisper-PMFA
Building on the successful use of pre-trained ASR models in speaker verification (Zhang et al., 2022; Cai et al., 2023), Zhao et al. (2024) recently proposed Whisper-PMFA (Partial Multi-Scale Feature Aggregation using Whisper) to leverage the capabilities of Whisper, a large-scale multilingual ASR model based on transformer architecture. Whisper-PMFA adapts Whisper for speaker verification by selectively concatenating frame-level outputs from specific transformer layers rather than aggregating features from all layers. This approach not only reduces computational overhead but also enhances performance by minimizing the integration of irrelevant information from lower-impact layers.
Inspired by this, we studied the performance of Whisper-PMFA in this challenge. Since Whisper was not trained for the speaker recognition task, we applied both training stages to Whisper-PMFA. Additionally, before the full training stage, we froze the Whisper parameters and fine-tuned the model for five epochs to prevent updating the pre-trained model in the wrong direction due to the random initialization of newly added components. We submitted two Whisper-PMFA-based systems (S4 and S5) to this challenge, differing only in the AAM-Softmax margin used during the domain adaptation phase: 0.35 for S4 and 0.2 for S5.
@@ 4.4 Feature Extraction
80-dimensional log Mel filter bank energies with a 25ms window and 10ms frame-shift were extracted for our ResNet-based models. Voice activity detection (VAD) was not applied, and all features were mean-normalized. Likewise, 80-dimensional log magnitude Mel spectrograms consistent with the pre-trained Whisper were utilized for training Whisper-PMFA.
@@ 4.5 Backend
Afterward, AS-Norm (Wang et al., 2020) was used for score normalization, using 1,620 cohorts obtained from speaker-wise averaging of all embeddings in the training subset of the challenge dataset. The top 300 most similar scores were selected to compute the mean and standard deviation for normalization.
@@
Methods Development MinDCF0.01 EER(%) Whisper-PMFA (T1) 0.0234 0.9253 + Domain adaptation (T2) 0.0177 0.6273 ++ AS-Norm 0.0161 0.6126
Table 4: Ablation study on Whisper-PMFA.
@@ 5 Results
Table 2 shows the evaluation results of our single and fusion systems on the development and evaluation subsets of the challenge after applying AS-Norm and rejecting TW trials. The results indicate that the Whisper-PMFA method outperforms the widely used ResNet architecture with random initialization, conforming to the findings of previous studies on the effectiveness of adapting pre-trained ASR models for speaker verification. However, it can be observed from the results that the ResNets pre-trained on approximately twice the data (VoxCeleb 1&2) can considerably surpass Whisper-PMFA after a well-designed domain adaptation stage, which highlights the importance of large-scale pre-training in improving the generalization ability of speaker verification models.
In addition, Figure 1 presents the Detection Error Tradeoff (DET) curves of the best-performing system for different categories of evaluation data. The results indicate that the model generally performs better on Persian phrases, which is expected given that the DeepMine dataset was collected from native Persian speakers, many of whom are likely less fluent in English. Furthermore, the results show noticeably higher performance for male speakers compared to female speakers. This disparity is not solely due to the inherent challenges of verifying female voices, but is also influenced by the specific characteristics of the DeepMine dataset, as discussed in its original description (Zeinali et al., 2018, 2019a) and in the official challenge results paper (Zeinali et al., 2025).
Figure 1: DET curves of our best-performing system.
@@ 5.1 Ablation Study
5.2 Comparison with Other Teams
To contextualize our performance, we report in Table 5 the evaluation results of our best system alongside the top-performing submissions in Task 1 of the international division of the TdSV Challenge. Team names and scores are taken directly from the official challenge results paper (Zeinali et al., 2025), which also provides brief descriptions and comparisons of the proposed architectures. As shown, our system achieves a lower MinDCF than the team ranked third in the international division.
Team MinDCF0.01 EER(%) Team 04 (Sreekanth, 2024) 0.0297 1.132 Team 08 0.0326 1.013 Our System 0.0358 1.246 Team 02 0.0379 1.164 Team 01 0.0504 2.245
Table 5: Evaluation results for our best system and the top-ranked teams in Task 1 of the international division of TdSV.
@@ 6 Conclusion
In this paper, we present our system for Task 1 of the Iranian division of the Text-dependent Speaker Verification (TdSV) Challenge 2024, focusing on resource-constrained training for TdSV systems. Unlike previous methods that jointly model speaker-related and linguistic features, our approach leverages two independent pre-trained models for phrase and speaker verification. This design reduces the computational costs associated with joint modeling during training while fully utilizing the capabilities of pre-trained models to achieve competitive performance. Our best system achieved a MinDCF of 0.0358 on the evaluation subset, securing first place in the challenge.